Session: Nov 1997 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (Tony Foley at Sun Nov 16 10:58:36 MET 1997) Failure
No phasecal. Scans from 1400-1530 seem OK, but akk dara from 1824-1954 is lost because of tape problems. The system reported vacuum loss, although it is not clear whether this was true. A second tape was used (OVLB0232) although no useful data was recorded on it.

Jodrell (Alastair Gunn at Sun Nov 16 15:11:00 MET 1997) Minor failures
FS crash at about 14:22, restarted at 15:17. Track 21 dead. Some illegal mode error messages. No other known problems.

Torun (K.Borkowski at Sun Nov 16 19:24:44 MET 1997) Failure
A few hours before this observation the controller of one of the motors on the elevation axis of our telescope went wrong.

Onsala (A. Polatidis at Tue Nov 18 13:35:55 MET 1997) Minor failures
Lost the first two scans because we could not establish vacuum. Changed the tape and everything was OK after this. Signal only on VC01 and VC03 (LCP).

Noto (C.Trigilio at Tue Nov 18 13:39:11 MET 1997) Success

Effelsberg (dgraham at Thu Nov 20 16:40:46 MET 1997) Minor failures
high Tsys between 1600 and 1700, presumably caused by rain

Medicina (dallacasa at Fri Nov 21 16:14:17 MET 1997) Success
No problems encountered

Calibration (Fredrik T Rantakyro at Tue Dec 9 15:03:55 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name vt907.antab Read the comments in the file to get further information.